OpenStudy (anonymous):

Find the 6th term for the sequence with the given formula: tn = 1 + 5n 11 30 6 31

OpenStudy (michele_laino):

If I replace \(n=1\) I get: \(t_1=1+5 \cdot 1=1+5=6\) so, please replace \(n=6\) using the same procedure

OpenStudy (michele_laino):

hint: \(t_6=1+5 \cdot 6=1+30=...?\)